A tungsten wire has a radius of 0.072 mm and is heated from 20.0 to 1410 °C. The temperature coefficient of resistivity is α = 4.5 10-3 (C°)−1. When 240 V is applied across the ends of the hot wire, a current of 1.9 A is produced. How long is the wire? Neglect any effects due to the thermal expansion of the wire.

_________________m
Answer #1

A=T8² - T (0.072 x 10 3 2 - 1.62 x 108 m² from ohns low R-V = 240 - 126.0 R = R. (1+x (1410-20). 126 = R. (1+ 4.500-3 (1390)